OpenStudy (anonymous):

9p-15=p+5p-2p linear equation

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

@Leong its the same thing

OpenStudy (anonymous):

What are you looking for?

OpenStudy (anonymous):

p

OpenStudy (anonymous):

@Mikeyy1992 ?

OpenStudy (sepeario):

Add the terms together?

OpenStudy (anonymous):

ok

OpenStudy (sepeario):

9p-15=4p 5p=15 p=?

OpenStudy (anonymous):

3 i allready know the answer i need help ith the steps

OpenStudy (anonymous):

@mathstudent55 @peachpi

OpenStudy (anonymous):

@sepeario

OpenStudy (anonymous):

@satellite73

OpenStudy (mathstudent55):

Step 1. Combine all like terms where they are. In your case, combine all like terms on the right side \(9p-15=p+5p-2p\) \(9p-15=4p\) There are no like terms on the left side, so step 1 is done. Ok so far?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Step 2. Now we want all terms with p on the left side, so we subtract 4p from both sides. \(~~~9p-15=~~~4p\) \(-4p ~~~~~~~~~~~~~-4p\) We get: \(5p - 15 = 0\)

OpenStudy (mathstudent55):

Step 3. Now we want the number (with no variable) on the right side, so we add 15 to both sides. \(~~~5p - 15 ~~~= ~~~~~~0\) \(~~~~~~~~+15~~~~~~~+15\) We get: \(5p = 15\) Ok so far?

OpenStudy (anonymous):

yes then divide then 3 THANKS. ;)

OpenStudy (mathstudent55):

Correct. Step 4. We want p alone, so we divide both sides by 5 \(\dfrac{5p}{5} = \dfrac{15}{5} \) \(p = 3\)